OpenStudy (anonymous):

the zeroes of the pol.p(x)=(x+5)(x+7) are

Parth (parthkohli):

The values that evaluate it to zero. What do you get?

Parth (parthkohli):

Hint: Some number multiplied by 0 is 0.

OpenStudy (anonymous):

whts d answer :P

Parth (parthkohli):

There are 2 possibilities here. Either x + 5 = 0 -OR- x + 7= 0 Solve them individually and they will be the values of x.

OpenStudy (anonymous):

solve these: x+5=0 x+7=0

OpenStudy (apoorvk):

For a polynomial of the form (x+a)(x+b), zeroes are "-a" and "-b" - can you guess why?? ;)

OpenStudy (anonymous):

nooooooooooooooo

OpenStudy (anonymous):

week in maths hate maths............

Parth (parthkohli):

Haha apoorv is correct. Try to substitute a and b with some numbers.

OpenStudy (apoorvk):

Okay, substitute '-a' into that equation - and you 'll know - think think!

OpenStudy (anonymous):

can any1 giv d answer

OpenStudy (anonymous):

the answer is what makes the statement (x+5)(+7)=0 true.

OpenStudy (apoorvk):

Not talking about hit and trial - @ParthKohli And Uh-oh, @Parvathysubhash , no one can just give you the answer - though we 'll help you in finding it :)

Parth (parthkohli):

x + 5 = 0 OR x + 7 = 0 Seriously, you just have to do one step in both. Just subtract both sides by 5 and 7 respectively.

OpenStudy (apoorvk):

You just gotta find that no. which reduces this polynomial to 'zero' - now if either of (x+5) or (x+7) go zero, the whole thing becomes zero as well, right? so, when can (x+5) be zero? or for that matter (x+7) ? - what Parth says above^.

OpenStudy (anonymous):

can u shw d steps full step so i can easly copy to ma note book

OpenStudy (apoorvk):

Lol No :D

OpenStudy (anonymous):

plzzzzzzzzzzzzzzzzzzzzzzzzzz

OpenStudy (apoorvk):

Follow us - it's so stupidly easy - you're gonna laugh at this once it's done!

OpenStudy (apoorvk):

No plzzzzzzzzzzzzzzzzzz :D

OpenStudy (anonymous):

no problem i hav to submit 2marrow no time pllzzzzzzzzzzzzz

OpenStudy (apoorvk):

Okay just tell me, x+7=0 how do you solve this?

OpenStudy (anonymous):

we will get -7

OpenStudy (anonymous):

rhyt

OpenStudy (apoorvk):

yeah!! so that was easy isn't it? same for x+5=0? what do you think?

OpenStudy (anonymous):

-7 and -5 we will gt

OpenStudy (apoorvk):

yeah right-o!!

OpenStudy (apoorvk):

so now can you understand what I was trying to say in my first reply up there^?

OpenStudy (anonymous):

noooo

OpenStudy (apoorvk):

The values of 'x' which reduce a polynomial to zero are called its roots or zeroes. Now, when I put in x=-a, in some polynomial (x+a)(x+b), what happens? (Don't fret this is easy. Forget everything else, Just relax and substitute '-a' instead of 'x'.)

OpenStudy (anonymous):

can u show it

OpenStudy (apoorvk):

Do you want the presentation of the problem?

OpenStudy (anonymous):

i need d ful steps

OpenStudy (apoorvk):

Well. If it helps you understand then... here we go. (x+5)(x+7) = 0 so x+5=0 and x+7 =0 or, x= -5 and -7. You couldn't do this much yourself?? :O

OpenStudy (anonymous):

dis i knw only dis much steps only dt i ma doubt dude....

OpenStudy (apoorvk):

Lol, hope doubts in this are solved.

OpenStudy (anonymous):

ysss